# Seek in the Dark: Reasoning via Test-Time Instance-Level Policy Gradient in Latent Space

### v1

@@ A.2 Theoretical Analysis: Independent Updating
Remark A.14.
[…]
Based on this theorem, we can easily derive Corollary A.11.
Corollary A.15.
NP $\subset$ NEXP $=$ MIP-Bounded
Appendix B Methods of TTIA
We list the formulations of two classical TTA methods:
    $x^{*}=\operatorname*{arg\,max}_{x}\pi(x|c)$ (11)
  • Best-of-N (BoN): Given $N$ i.i.d sequences $x_{(1)},x_{(2)},\ldots,x_{(N)}\sim\pi(\cdot|c)$,
    $x^{*}=\operatorname*{arg\,max}_{x\in\{x_{(1)},x_{(2)},\ldots,x_{(N)}\}}R(x,c)$ (12)
Appendix C Experimental Details
@@ C.1 Prompt Designation
The prompts are listed in Tables 18, 16, 17 and 15.
C.2 Backbone
@@ C.3 Baselines
  • Chain-of-Thought (CoT): CoT refers to a structured, sequential approach to problem-solving, wherein complex tasks are decomposed into intermediate steps, each explicitly articulated to facilitate logical progression toward a solution. In our experiments, CoT follows the prompt listed in Section C.1.
@@
  • Best-of-N (BoN): This method involves generating $N$ independent samples or candidates from a given model, evaluating each based on a predefined objective function, and selecting the candidate that maximizes the desired reward. In our experiment, since the average update iterations of LatentSeek is less than 2, we select $N=3$, which is larger, for BoN. The reward function for BoN remains the same as that of LatentSeek, which is performed under self-reward mechanism.
@@
[…]
@@ C.4 GSM8K
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table 9 and Table 10, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The detailed specifications of these four evaluation prompts are provided in Tables 15, 18, 16 and 17.The prompt structure is consistently applied across all backbones and both prompts. Notably, for Llama-3.1-8B-Instruct, Qwen2.5-1.5B-Instruct and Qwen2-7B-Instruct in LatentSeek (Self) experiment using prompt 1, a new format reward following the methodology of DeepSeek-R1 is incorporated, with weights of 3, 2, and 2 assigned to format-based criteria, respectively.
@@
We offer more scaling figures in Figure 3.
@@
(f) Mistral-7B-Instruct-v0.3
@@
LatentSeek (Self) Qwen2.5-1.5B-Instruct 1024 1 1 3090 10 Adam 0.2 bf16 10
LatentSeek (Self) Qwen2.5-1.5B-Instruct 1024 2 1 3090 0.03 Adam 0.2 bf16 10
@@
LatentSeek (Self) Mistral-7B-Instruct-v0.3 1024 1 1 L40 0.05 Adam 0.2 bf16 10
LatentSeek (Self) Mistral-7B-Instruct-v0.3 1024 2 1 L40 0.05 Adam 0.2 bf16 10
@@
LatentSeek (PSRM) Mistral-7B-Instruct-v0.3 1024 1 1 A100 0.05 Adam 0.2 bf16 10
LatentSeek (PSRM) Mistral-7B-Instruct-v0.3 1024 2 1 A100 0.05 Adam 0.2 bf16 10
C.5 MATH-500
@@
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table 11 and Table 12, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The prompt structure is consistently applied across all backbones and both prompts. Notably, for Qwen2-7B-Instruct and Llama-3.1-8B-Instruct in LatentSeek (Self) experiment using prompt 1, the weight ratios are adjusted to 1:1:1:2. Additionally, for Qwen2-7B-Instruct, Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct in LatentSeek (Self) experiment using prompt 1, a new format reward following the methodology of DeepSeek-R1 is incorporated, with weight of 2 assigned to format-based criteria.
@@
LatentSeek (Self) Qwen2.5-1.5B-Instruct 1024 2 1 A100 0.03 Adam 0.1 bf16 10
@@
LatentSeek (Self) Mistral-7B-Instruct-v0.3 1024 1 1 3090 0.03 Adam 0.1 bf16 10
LatentSeek (Self) Mistral-7B-Instruct-v0.3 1024 2 1 3090 0.03 Adam 0.1 bf16 10
@@
LatentSeek (PSRM) Mistral-7B-Instruct-v0.3 1024 1 1 3090 0.03 Adam 0.1 bf16 10
LatentSeek (PSRM) Mistral-7B-Instruct-v0.3 1024 2 1 3090 0.03 Adam 0.1 bf16 10
C.6 AIME2024
@@
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table 13 and Table 14, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The prompt structure is consistently applied across all backbones and both prompts.
@@
LatentSeek (Self) Mistral-7B-Instruct-v0.3 1024 1 1 3090 0.03 Adam 0.1 bf16 10
LatentSeek (Self) Mistral-7B-Instruct-v0.3 1024 2 1 3090 0.05 Adam 0.2 bf16 10
@@
LatentSeek (PSRM) Mistral-7B-Instruct-v0.3 1024 1 1 3090 0.03 Adam 0.1 bf16 10
LatentSeek (PSRM) Mistral-7B-Instruct-v0.3 1024 2 1 3090 0.03 Adam 0.1 bf16 10
@@
Appendix D Qualitative Analysis and Case Studies

### v2

@@ C.2 Theoretical Analysis: Independent Updating
Remark C.14.
NEXP is the complexity class that a non-deterministic Turing Machine can solve in exponential time, which means that $\text{NP}\subset\text{NEXP}$.
Based on this theorem, we can easily derive Corollary˜C.11.
Corollary C.15.
NP $\subset$ NEXP == MIP-Bounded
Appendix D Derivation of Policy Gradient
Our target is to derive Equation˜7:
$\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})]_{t}=\mathbb{E}_{\mathbf{x}\sim\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})}\left[R(\mathbf{x},\mathbf{c})\nabla_{z_{t}}\log\pi(x_{t}\mid z_{t})\right]$
Starting from the objective:
$\mathcal{J}(\mathbf{z}):=\mathbb{E}_{\mathbf{x}\sim\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})}\left[R(\mathbf{x},\mathbf{c})\right]$
By taking gradient of $\mathbf{z}$, we have:
$\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})=\nabla_{\mathbf{z}}\int_{\mathbf{x}}R(\mathbf{x},\mathbf{c})\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})d\mathbf{x}=\int_{\mathbf{x}}R(\mathbf{x},\mathbf{c})\nabla_{\mathbf{z}}\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})d\mathbf{x}$
Leveraging $\nabla_{\mathbf{z}}\log\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})=\frac{1}{\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})}\nabla_{\mathbf{z}}\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})$:
$\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})=\int_{\mathbf{x}}R(\mathbf{x},\mathbf{c})\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})\nabla_{\mathbf{z}}\log\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})d\mathbf{x}$
According to $\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})=\prod\limits_{t=1}^{N}\pi(x_{t}\mid z_{t})\prod\limits_{t=N+1}^{T}\pi(x_{t}|x_{<t},\mathbf{c})$, we have:
$\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})=\int_{\mathbf{x}}R(\mathbf{x},\mathbf{c})\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})\nabla_{\mathbf{z}}(\sum\limits_{t=1}^{N}\log\pi(x_{t}\mid z_{t})+\sum_{t=N+1}^{T}\log\pi(x_{t}|x_{<t}))d\mathbf{x}$
As the second term doesn’t have $\mathbf{z}$, thus we have:
$\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})=\int_{\mathbf{x}}R(\mathbf{x},\mathbf{c})\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})\nabla_{\mathbf{z}}(\sum\limits_{t=1}^{N}\log\pi(x_{t}\mid z_{t}))d\mathbf{x}=\mathbb{E}_{\mathbf{x}\sim\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})}[R(\mathbf{x},\mathbf{c})\nabla_{\mathbf{z}}(\sum_{t=1}^{N}\log\pi(x_{t}|z_{t}))]$
Therefore, the $t$-th term is:
$\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})]_{t}=\mathbb{E}_{\mathbf{x}\sim\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})}[R(\mathbf{x},\mathbf{c})\nabla_{z_{t}}\log\pi(x_{t}|z_{t})],$
which is exactly what we want.
Appendix E Additional Experimental Results
E.1 LatentSeek vs BoN (N=10)
In Table˜1, we show that Best-of-N (BoN) is outperformed by LatentSeek, even though BoN consumes a comparable compute consumption (See Figure˜4). One might also ask how LatentSeek compares to BoN when $N=10$; in this setting, BoN requires at least five times more sequence-level computation than LatentSeek. As reported in Table˜6, LatentSeek remains superior on all datasets and all backbone models.
Table 6: Additional Comparison Between LatentSeek and BoN (N=10) with Prompt 2.
Methods Model Qwen2 Qwen2.5 LLaMA3.1 Avg
7B 1.5B 7B 14B 8B
GSM8K
BoN 72.10 7.73 80.06 92.19 81.58 66.73
LatentSeek 80.21+8.11 44.20+36.47 85.06+5.00 92.72+0.53 83.70+2.12 77.18+10.45
MATH-500
BoN 42.40 28.80 56.40 63.20 45.60 47.28
LatentSeek 44.80+2.40 32.20+3.40 57.60+1.20 71.00+7.80 47.00+1.40 50.52+3.24
AIME2024
BoN 3.33 0.00 6.67 10.00 3.33 4.66
LatentSeek 3.33+0.00 3.33+3.33 13.33+6.66 10.00+0.00 6.67+3.33 7.33+2.67
E.2 Qwen3 AIME Performance
Figure 3: Performance comparison of the Qwen3-4B-Instruct model on the AIME2024 dataset. The chart illustrates the accuracy of LatentSeek against the Chain-of-Thought (CoT) and Best-of-N (BoN) baselines across two distinct prompt formats.
This section provides a detailed study on the performance of the Qwen3-4B-Instruct-2507 model on the challenging AIME2024 dataset. Figure˜3 illustrates the accuracy of LatentSeek compared against the Chain-of-Thought (CoT) and Best-of-N (BoN) baselines under two prompt formats.
As shown in the figure, under Prompt 1, while the BoN strategy fails to improve upon the CoT baseline (both at 63.3% accuracy score), LatentSeek delivers a significant 10 points uplift to 73.3%. This advantage holds under Prompt 2, where our method’s accuracy of 56.7% surpasses CoT by 16.7 points and BoN by 3.4 points. Taken together, these findings demonstrate that LatentSeek consistently improves upon strong baselines and is valuable in scenarios where conventional search methods are insufficient.
E.3 Token Efficiency Comparison with BoN Baseline
This section provides a quantitative analysis of the token efficiency of LatentSeek, comparing it against the Best-of-N (BoN) baseline reported in Table˜1. The analysis measures the average number of tokens consumed per problem instance under the JSON prompt format (Prompt 2) on the GSM8K and MATH-500 datasets.
As shown in Figure˜4, LatentSeek achieves higher token efficiency than BoN in all of cases across different model families and sizes, while attaining superior reasoning accuracy as reported in Table˜1. These results demonstrates that LatentSeek can enhance performance without incurring additional token overhead for all models, highlighting the efficiency of latent space optimization.
(a) GSM8K dataset.
(b) MATH-500 dataset.
Figure 4: Comparison of average tokens consumed per problem between LATENTSEEK and the Best-of-N (BoN) baseline when using Prompt 2. The token consumption is comparable for the two methods, with BoN consuming slightly more tokens in total. For experimental results where BoN consumes extremely more calculation, please refer to Table˜6.
E.4 Greedy Decoding vs Sampling
Our experiments show that greedy decoding closely approximates categorical sampling while being a better choice for reducing stochastic error. As detailed in Table˜7, greedy decoding performs slightly better than the sampling method. We hypothesize that this is because sampling can amplify the approximation error when replacing an expected mean with an empirical one. Therefore, we use greedy decoding for our main experiments.
Table 7: LatentSeek: Greedy Decoding vs Sampling (Temperature 0.7) on the MATH-500 dataset.
Methods Model Qwen2 Qwen2.5 LLaMA3.1
7B 1.5B 7B 14B 8B
Prompt 1
Sampling 56.80 50.00 74.00 74.00 49.60
Greedy Decoding 57.40 55.60 75.60 80.00 54.60
Prompt 2
Sampling 44.40 22.40 55.80 68.40 45.40
Greedy Decoding 44.80 34.40 57.60 71.00 47.00
E.5 Limitations of Outcome Reward Model
The performance is described in Liu et al. (2024b), with results presented in Table˜8. Although this model achieves an average score of 64.79—representing an improvement of 3.23 points over the Chain-of-Thought (CoT) method—it remains significantly inferior to the performance attained using a self-rewarding mechanism, which reaches 77.18. These results highlight the limitations of current outcome-based reward models in scenarios lacking ground-truth data, underscoring the need for the development of more robust and effective outcome reward models.
Methods Model Qwen2 Qwen2.5 LLaMA3.1 Avg
7B 1.5B 7B 14B 8B
GSM8k: Prompt 2
CoT 65.20 15.31 66.41 91.81 69.07 61.56
LatentSeek (Math) 61.18 28.28 66.64 91.05 76.80 64.79
LatentSeek (Self) 80.21 44.20 85.06 92.72 83.70 77.18
E.6 Performance vs Fraction Ratio
Figure˜5 illustrates the performance trends of various instruction-tuned language models across different fraction ratios, which range from 0.1 to 0.8. Here, we also experimented with another model: Mistral-7B-Instruct (Jiang et al., 2023). Performance is reported as a percentage and plotted on the y-axis, while the fraction ratio is shown on the x-axis.
Figure 5: Performance vs Fraction Ratio
Appendix F Experimental Details
@@ F.1 Prompt Designation
The prompts are listed in Tables˜21, 19, 20 and 18.
F.2 Backbone
@@ F.3 Baselines
  • Chain-of-Thought (CoT): CoT refers to a structured, sequential approach to problem-solving, wherein complex tasks are decomposed into intermediate steps, each explicitly articulated to facilitate logical progression toward a solution. In our experiments, CoT follows the prompt listed in Section˜F.1.
@@
  • Best-of-NN (BoN): We draw $N$ independent candidates from the model, score each with a predefined objective, and select the candidate with the highest reward. In our experiments, we set $N=3$ so that the total computation is comparable to—while exceeding—that of LatentSeek. The BoN reward is computed under the same self-reward mechanism as that of LatentSeek.
  • Self-Consistency: We draw $N$ independent candidates from the model, and select the candidate that appears with the highest frequency. In our experiments, we set $N=3$ so that the total computation is comparable to—while exceeding—that of LatentSeek and align with that of BoN.
@@
$\backslash${"role": "system", "content": "You are a precise math question solver. Solve this math problem." $\backslash$},
@@
$\backslash${"role": "system", "content": "You are a precise math question solver. Solve this math problem." $\backslash$},
@@ F.4 GSM8K
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table˜12 and Table˜13, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The detailed specifications of these four evaluation prompts are provided in Tables˜18, 21, 19 and 20.The prompt structure is consistently applied across all backbones and both prompts. Notably, for Llama-3.1-8B-Instruct, Qwen2.5-1.5B-Instruct and Qwen2-7B-Instruct in LatentSeek (Self) experiment using Prompt 1, a new format reward following the methodology of DeepSeek-R1 is incorporated, with weights of 3, 2, and 2 assigned to format-based criteria, respectively.
@@
We offer more scaling figures in Figure˜6.
@@
Figure 6: GSM8kCobbe et al. (2021) Prompt 2 Accuracy changes with respect to the increasing number of iterations. Orange: Perfect Sharp Reward Model. Blue: Self Reward Model.
@@
LatentSeek (Self) Qwen2.5-1.5B-Instruct 1024 1 1 3090 0.03 Adam 0.05 bf16 10
LatentSeek (Self) Qwen2.5-1.5B-Instruct 1024 2 1 3090 10 Adam 0.2 bf16 10
@@
F.5 MATH-500
@@
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table˜14 and Table˜15, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The prompt structure is consistently applied across all backbones and both prompts. Notably, for Qwen2-7B-Instruct and Llama-3.1-8B-Instruct in LatentSeek (Self) experiment using prompt 1, the weight ratios are adjusted to 1:1:1:2. Additionally, for Qwen2-7B-Instruct, Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct in LatentSeek (Self) experiment using Prompt 1, a new format reward following the methodology of DeepSeek-R1 is incorporated, with weight of 2 assigned to format-based criteria.
@@
LatentSeek (Self) Qwen2.5-1.5B-Instruct 1024 2 1 3090 0.03 Adam 0.05 bf16 10
@@
F.6 AIME2024
@@
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table˜16 and Table˜17, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The prompt structure is consistently applied across all backbones and both prompts.
@@
Table 22: Estimated FLOPs for a single forward pass of the LLaMA3.1-8B backbone model with a context length of 512 tokens. The total is derived by summing the operational costs for both the prefill and decode stages.
Prefill Stage (512 tokens) Decode Stage (512 tokens)
Component FLOPs Component FLOPs
q_proj 17.2G q_proj 33.6M
k_proj 4.3G k_proj 8.4M
v_proj 4.3G v_proj 8.4M
out_proj 17.2G out_proj 33.6M
gate_proj 60.1G gate_proj 117M
up_proj 60.1G up_proj 117M
down_proj 60.1G down_proj 117M
qk_matmul 2.1G qk_matmul 4.2M
sv_matmul 2.1G sv_matmul 4.2M
softmax 41.9M softmax 81.9K
attn_norm 14.7M attn_norm 28.7K
mlp_norm 14.7M mlp_norm 28.7K
attn_add 2.1M attn_add 4.1K
mlp_add 2.1M mlp_add 4.1K
mlp_act 4.2M mlp_act 8.2K
lm_head 525M lm_head 525M
Appendix G Detailed FLOPs Calculation
This section provides a detailed breakdown of the Floating Point Operations (FLOPs) required for the Genius baseline and our proposed LatentSeek method, using LLaMA3.1-8B-Instruct as the backbone model.
G.1 Forward Pass FLOPs Estimation
The total FLOPs for a single forward pass is estimated by summing the operational costs of its constituent components, as detailed in Table˜22. The total FLOPs for one forward pass is calculated as follows:
$\text{FLOPs}_{\text{forward}}$ $=(227.5\times 10^{9})+(1573.1\times 10^{6})+(155.7\times 10^{3})$
$\approx 2.29\times 10^{11}\text{ FLOPs }$
G.2 Total FLOPs for Genius
On the GSM8K dataset, the Genius framework involoves a full training phase on 100000 instances followed by testing on 1319 instances. We assume a backward pass is approximately twice as computationally expensive as a forward pass.
  • Training FLOPs: For each of the 100000 instances, one forward pass and one full backward pass are required.
    $\text{FLOPs}_{\text{train}}=100000\times(\text{FLOPs}_{\text{forward}}+2\times\text{FLOPs}_{\text{forward}})=300000\times\text{FLOPs}_{\text{forward}}$
  • Testing FLOPs: For each of the 1319 instances, one forward pass is required.
    $\text{FLOPs}_{\text{test}}=1319\times\text{FLOPs}_{\text{forward}}$
$\textbf{FLOPs}_{\text{Genius}}$ $=\text{FLOPs}_{\text{train}}+\text{FLOPs}_{\text{test}}$
$=(300000+1319)\times(2.29\times 10^{11})$
$\approx 6.90\times 10^{16}\text{FLOPs}$
G.3 Total FLOPs for LatentSeek
LatentSeek operates exclusively at test time on 1319 instances. Each iteration involoves two forward passes (one for generation, one for reward evaluation) and one partial backward pass through only the LM head. From Table˜4, the average number of iterations on GSM8K is 1.27.
  • Number of Forward Passes: $1319\text{ instances}\times 1.27\text{ iter/inst}\times 2\text{ forward/iter}\approx 3350$
  • Number of Backward Passes: $1319\text{ instances}\times 1.27\text{ iter/inst}\times 1\text{ forward/iter}\approx 1675$
  • Cost of Backward Pass: The backward pass is only through the LM head (525 MFLOPs forward).
[…]
The total extimated FLOPs for LatentSeek is:
$\textbf{FLOPs}_{\text{LATENTSEEK}}$ $=(3350\times\text{FLOPs}_{\text{forward}})+(1675\times\text{FLOPs}_{\text{backward\_LATENTSEEK}}$
$=(3350\times 2.29\times 10^{11})+(1675\times 1.05\times 10^{9})$
$\approx 7.67\times 10^{14}+1.76\times 10^{12}$
$\approx 7.69\times 10^{14}\text{FLOPs}$
G.4 Efficiency Threshold Analysis
We calculate the number of inference instances ($x$) at which the total computational cost of using the pre-trained Genius framework would equal the cost of using LatentSeek. This threshold is found by solving the following equation, where the left side represents the total cost of Genius and the right side represents the total cost of LatentSeek for $x$ inferences:
$6.90\times 10^{16}+x\cdot(2.29\times 10^{11})=1.27\times(2\times 2.29\times 10^{11}+1.05\times 10^{9})\cdot x$
Solving for $x$ yields the threshold:
$x\approx 1.94\times 10^{5}$
This calculation demonstrates that LatentSeek remains more efficient than the Genius method up to approximately $1.94\times 10^{5}$ inference instances, which exceeds the size of Genius’s training set ($1.00\times 10^{5}$ examples).
Appendix H Qualitative Analysis and Case Studies